# The computational power of discrete chemical reaction networks with bounded executions

### v2

@@ Abstract
Chemical reaction networks (CRNs) model systems where molecules interact according to a finite set of reactions such as $A+B\to C$, representing that if a molecule of $A$ and $B$ collide, they disappear and a molecule of $C$ is produced. CRNs can compute Boolean-valued predicates $\phi:\mathbb{N}^{d}\to\{0,1\}$ and integer-valued functions $f:\mathbb{N}^{d}\to\mathbb{N}$; for instance $X_{1}+X_{2}\to Y$ computes the function $\min(x_{1},x_{2}).$
@@
However, if no initial leader is allowed, all species vote, and the CRN is “noncollapsing” (does not shrink from initially large to final $O(1)$ size configurations), then execution bounded CRNs are severely limited, able to compute only eventually constant predicates. A key tool is to characterize execution bounded CRNs as precisely those with a nonnegative linear potential function that is strictly decreased by every reaction, a result that may be of independent interest.
@@ 1 Introduction
Chemical reaction networks (CRNs) are a fundamental tool for understanding and designing molecular systems. By abstracting chemical reactions into a set of finite, rule-based transformations, CRNs allow us to model the behavior of complex systems. For instance, the CRN with a single reaction $2X\rightarrow Y$, produces one $Y$ every time two $X$ molecules randomly react together, effectively calculating the function $f(x)=\lfloor x/2\rfloor$ if the initial count of $X$ molecules is interpreted as the input and $Y$ as the output. A commonly studied special case of CRNs is the population protocol model of distributed computing [Angluin2006ComputationalPower], in which each reaction has exactly two reactants and two products, e.g., $A+B\to C+D$. This model assumes idealized conditions where reactions can proceed indefinitely, constrained only by the availability of reactants in the well-mixed solution.
Precisely the semilinear predicates $\phi:\mathbb{N}^{d}\to\{0,1\}$ and functions $f:\mathbb{N}^{d}\to\mathbb{N}$ can be computed stably, roughly meaning that the output is correct no matter the order in which reactions happen. In population protocols or other CRNs with a finite reachable configuration space, this means that the output is correct with probability 1 under a stochastic scheduler that picks the next molecules to react at random. However, existing constructions to compute semilinear predicates and functions use CRNs with unbounded executions, meaning that it is possible to execute infinitely many reactions from the initial configuration. CRNs with bounded executions have several advantages. With an absolute guarantee on how many reactions will happen before the CRN terminates, wet-lab implementations need only supply a bounded amount of fuel to power the reactions. Such CRNs are simpler to reason about: each reaction brings it “closer” to the answer. They also lead to a simpler definition of stable computation than is typically employed: an execution bounded CRN stably computes a predicate/function if it gets the correct answer after sufficiently many reactions.
To study this topic, we limit the classical, discrete CRN model to networks that must eventually reach a configuration where no further reactions can occur, regardless of the sequence of reactions executed. By guaranteeing a finite endpoint for CRN computations and later integrating the concept of decreasing potential, we aim to align our models more closely with their implementations in the physical world.
This restriction is nontrivial because the techniques in [Chen2012DeterministicFunction] and [Doty2013Leaderless] rely on reversible reactions catalyzed by species we expect to be depleted once a computational step has terminated. This trick seems to add computational power to our system by undoing certain reactions as long as a specific species is present. Consider the following CRN computing $f(x_{1},x_{2},x_{3})=\min\left(x_{1}-x_{2},x_{3}\right)$. The input values $x_{i}$ are given as counts of copies of $X_{i}$, and the count of $Z$ molecules in the stable output:
@@
Reactions (1) and (2) compute $x_{1}-x_{2}$, storing the result in the count of $Y$. Next, reaction (3) can be applied exactly $\min(y,x_{3})$ times. But since the order of reactions is a stochastic process, we might consume copies of $Y$ in (3), before all of $x_{2}$ is subtracted from it. Therefore, we add reaction (4), using $X_{2}$ as a catalyst to undo reaction (3) as long as copies of $X_{2}$ are present, indicating that the first step of computation has not terminated. A similar technique is used in [Chen2012DeterministicFunction], where semilinear sets are understood as a finite union of linear sets, shown to be computable in parallel by CRNs. A reversible, catalyzed reaction finally converts the output of one of the CRNs to the global output. Among other questions, we explore how the constructions of [Chen2012DeterministicFunction] and [Doty2013Leaderless] can be modified to provide equal computational power while guaranteeing bounded execution.
Section 3 defines execution boundedness (Definition 3.1). Furthermore, we introduce alternative characterizations of the class for use in later proofs, such as the lack of self-covering execution paths. Section 4 and 5 contain the main positive results of the paper and provide the concrete constructions used to decide semilinear sets and functions using execution bounded CRNs whose initial configurations contain a single leader. Section 6 discusses the limitations of execution bounded CRNs, introducing the concept of a “linear potential function” as a core characterization of these systems. We demonstrate that entirely execution bounded CRNs that are leaderless and non-collapsing (such as all population protocols), can only stably decide trivial semilinear predicates: the eventually constant predicates (Definition 6.11).
@@ 2 Preliminaries
We use established notation from [Chen2012DeterministicFunction, Doty2013Leaderless] and stable computation definitions from [Angluin2006ComputationalPower] for (discrete) chemical reaction networks.
@@ 2.1 Notation
Let $\mathbb{N}$ denote the nonnegative integers. For any finite set $\Lambda$, we write $\mathbb{N}^{\Lambda}$ to mean the set of functions $f:\Lambda\rightarrow\mathbb{N}$. Equivalently, $\mathbb{N}^{\Lambda}$ can be interpreted as the set of vectors indexed by the elements of $\Lambda$, and so $\mathbf{c}\in\mathbb{N}^{\Lambda}$ specifies nonnegative integer counts for all elements of $\Lambda$. $\mathbf{c}(i)$ denotes the $i$-th coordinate of $\mathbf{c}$, and if $\mathbf{c}$ is indexed by elements of $\Lambda$, then $\mathbf{c}(Y)$ denotes the count of species $Y\in\Lambda$.
@@ 2.2 Chemical Reaction Networks
A chemical reaction network (CRN) is a pair $\mathcal{C}=(\Lambda,R)$, where $\Lambda$ is a finite set of chemical species, and $R$ is a finite set of reactions over $\Lambda$, where each reaction is a pair $(\mathbf{r},\mathbf{p})\in\mathbb{N}^{\Lambda}\times\mathbb{N}^{\Lambda}$ indicating the reactants $\mathbf{r}$ and products $\mathbf{p}$. A population protocol [angluin2004computation] is a CRN in which all reactions $(\mathbf{r},\mathbf{p})$ obey $\|\mathbf{r}\|=\|\mathbf{p}\|=2$. We write reactions such as $A+2B\mathop{\rightarrow}\limits A+3C$ to represent the reaction ($(\{A,2B\},\{A,3C\})$. A configuration $\mathbf{c}\in\mathbb{N}^{\Lambda}$ of a CRN assigns integer counts to every species $S\in\Lambda$. When convenient, we use the notation $\left\{n_{1}S_{1},n_{2}S_{2},\ldots,n_{k}S_{k}\right\}$ to describe a configuration $\mathbf{c}$ with $n_{i}\in\mathbb{N}$ copies of species $S_{i}$, i.e., $\mathbf{c}(S_{i})=n_{i}$, and any species that is not listed is assumed to have a zero count. If some configuration $\mathbf{c}$ is understood from context, for a species $S$, we write #$\#S$ to denote $\mathbf{c}(S).$ A reaction $(\mathbf{r},\mathbf{p})$ is said to be applicable in configuration $\mathbf{c}$ if $\mathbf{r}\leqq\mathbf{c}$. If the reaction $(\mathbf{r},\mathbf{p})$ is applicable, applying it results in configuration $\mathbf{c}^{\prime}=\mathbf{c}-\mathbf{r}+\mathbf{p}$, and we write $\mathbf{c}\rightarrow\mathbf{c}^{\prime}$.
@@
Let $\mathbf{u}\in\mathbb{N}^{R}$. Then the vector $\mathbf{M}\mathbf{u}\in\mathbb{Z}^{\Lambda}$ represents the change in species counts that results from applying reactions by amounts described in $\mathbf{u}$. In the above example, if $\mathbf{u}=(2,1)$, then $\mathbf{M}\mathbf{u}=(-1,0,4)$, meaning that executing the first reaction twice ($\mathbf{u}_{1}=2$) and the second reaction once ($\mathbf{u}_{2}=1$) causes $S_{1}$ to decrease by 1, $S_{2}$ to stay the same, and $S_{3}$ to increase by 4.
@@ 2.3 Stable computation with CRNs
To capture the result of computations done by a CRN, we generalize the definitions to include information about how to interpret the final configuration after letting the CRN run until the result cannot change anymore (characterized below as stable computation). Computation primarily involves two classes of functions: 1. evaluating predicates to determine properties of the input (akin to deciding a set defined by these properties), and 2. executing general functions that map an input configuration to an output, denoted as $f:\mathbb{N}^{k}\rightarrow\mathbb{N}$.
@@
A chemical reaction decider (CRD) is a tuple $\mathcal{D}=(\Lambda,R,\Sigma,\Upsilon_{\mathrm{Y}},\Upsilon_{\mathrm{N}},\mathbf{s})$, where $(\Lambda,R)$ is a CRN, $\Sigma\subseteq\Lambda$ is the set of input species, $\Upsilon_{\mathrm{Y}}\subseteq\Lambda$ is the set of yes voters, and $\Upsilon_{\mathrm{N}}\subseteq\Lambda$ is the set of no voters. If $\Upsilon_{\mathrm{Y}}\cup\Upsilon_{\mathrm{N}}=\Lambda$, we say the CRD is all-voting. We define a global output partial function $\Phi:\mathbb{N}^{\Lambda}\longrightarrow\{0,1\}$ as follows. $\Phi(\mathbf{c})$ is undefined if either $\mathbf{c}=\mathbf{0}$, or if there exist $S_{\mathrm{N}}\in\Upsilon_{\mathrm{N}}$ and $S_{\mathrm{Y}}\in\Upsilon_{\mathrm{Y}}$ such that $\mathbf{c}\left(S_{\mathrm{N}}\right)>0$ and $\mathbf{c}\left(S_{\mathrm{Y}}\right)>0$. In other words, we require a unanimous vote as our output. We say $\mathbf{c}$ is stable if, for all $\mathbf{c}^{\prime}$ such that $\mathbf{c}\Rightarrow\mathbf{c}^{\prime}$, $\Phi(\mathbf{c})=\Phi(\mathbf{c}^{\prime}).$ We say a CRD $\mathcal{D}$ stably decides the predicate $\psi:\mathbb{N}^{\Sigma}\rightarrow\{0,1\}$ if, for any valid initial configuration $\mathbf{i}\in\mathbb{N}^{\Lambda}$ with $\mathbf{i}\upharpoonright\Sigma=\mathbf{i}_{0}$, for all configurations $\mathbf{c}\in\mathbb{N}^{\Lambda},\mathbf{i}\Rightarrow\mathbf{c}$ implies $\mathbf{c}\Rightarrow\mathbf{c}^{\prime}$ such that $\mathbf{c}^{\prime}$ is stable and $\Phi\left(\mathbf{c}^{\prime}\right)=\psi\left(\mathbf{i}_{0}\right)$. We associate to a predicate $\psi$ the set $A=\psi^{-1}(1)$ of inputs on which $\psi$ outputs 1, so we can equivalently say the CRD stably decides the set $A.$
@@
A chemical reaction computer $(CRC)$ is a tuple $\mathcal{C}=(\Lambda,R,\Sigma,Y,\mathbf{s})$, where $(\Lambda,R)$ is a CRN, $\Sigma\subset\Lambda$ is the set of input species, $Y\in\Lambda\backslash\Sigma$ is the output species, and $\mathbf{s}\in\mathbb{N}^{\Lambda\backslash\Sigma}$ is the initial context. A configuration $\mathbf{o}\in\mathbb{N}^{\Lambda}$ is stable if, for every $\mathbf{c}$ such that $\mathbf{o}\Rightarrow\mathbf{c},\mathbf{o}(Y)=\mathbf{c}(Y)$, i.e. the output can never change again. We say that $\mathcal{C}$ stably computes a function $f:\mathbb{N}^{k}\rightarrow\mathbb{N}$ if for any valid initial configuration $\mathbf{i}\in\mathbb{N}^{\Sigma}$ and any $\mathbf{c}\in\mathbb{N}^{\Lambda},\mathbf{i}\Rightarrow\mathbf{c}$ implies $\mathbf{c}\Rightarrow\mathbf{o}$ such that $\mathbf{o}$ is stable and $f(\mathbf{i}\upharpoonright\Sigma)=\mathbf{o}(Y)$ where $\mathbf{i}\upharpoonright\Sigma$ denotes restriction of $\mathbf{i}$ to $\Sigma$.
For a CRD or CRC with initial context $\mathbf{s}$ and input species $\Sigma$, we say a $\mathbf{i}$ is a valid initial configuration if $\mathbf{i}=\mathbf{s}+\mathbf{x}$, where $\mathbf{x}(S)=0$ for all $S\in\Lambda\setminus\Sigma$; i.e., $\mathbf{i}$ is the initial context plus only input species.table configuration.

### v3

@@ Abstract
Chemical reaction networks (CRNs) model systems where molecules interact according to a finite set of reactions such as $A+B\to C$, representing that if a molecule of $A$ and $B$ collide, they disappear and a molecule of $C$ is produced. CRNs can compute Boolean-valued predicates $\phi:\mathbb{N}^{d}\to\{0,1\}$ and integer-valued functions $f:\mathbb{N}^{d}\to\mathbb{N}$; for instance $X_{1}+X_{2}\to Y$ computes the function $\min(x_{1},x_{2}),$ since starting with $x_{i}$ copies of $X_{i}$, eventually $\min(x_{1},x_{2})$ copies of $Y$ are produced.
@@
However, if no initial leader is allowed, all species vote, and the CRN is “non-collapsing” (does not shrink from initially large to final $O(1)$ size configurations), then execution bounded CRNs are severely limited, able to compute only eventually constant predicates. A key tool is a characterization of execution bounded CRNs as precisely those with a nonnegative linear potential function that is strictly decreased by every reaction [6].
@@ 1 Introduction
Chemical reaction networks (CRNs) are a fundamental tool for understanding and designing molecular systems. By abstracting chemical reactions into a set of finite, rule-based transformations, CRNs allow us to model the behavior of complex chemical systems. For instance, the CRN with a single reaction $2X\rightarrow Y$, produces one $Y$ every time two $X$ molecules randomly react together, effectively calculating the function $f(x)=\lfloor x/2\rfloor$ if the initial count of $X$ is interpreted as the input and the eventual count of $Y$ as the output. A commonly studied special case of CRNs is the population protocol model of distributed computing [3], in which each reaction has exactly two reactants and two products, e.g., $A+B\to C+D$. This model assumes idealized conditions where reactions can proceed indefinitely, constrained only by the availability of reactants in the well-mixed solution.
Precisely the semilinear predicates $\phi:\mathbb{N}^{d}\to\{0,1\}$ [1] and functions $f:\mathbb{N}^{d}\to\mathbb{N}$ [5] can be computed stably, roughly meaning that the output is correct no matter the order in which reactions happen. In population protocols or other CRNs with a finite reachable configuration space, this means that the output is correct with probability 1 under a stochastic scheduler that picks the next molecules to react at random. However, existing constructions to compute semilinear predicates and functions use CRNs with unbounded executions, meaning that it is possible to execute infinitely many reactions from the initial configuration. CRNs with bounded executions have several advantages. With an absolute guarantee on how many reactions will happen before the CRN terminates, wet-lab implementations need only supply a bounded amount of fuel to power the reactions. Such CRNs are simpler to reason about: each reaction brings it “closer” to the answer. They also lead to a simpler definition of stable computation than is typically employed: an execution bounded CRN stably computes a predicate/function if it gets the correct answer after sufficiently many reactions.
To study this topic, we study networks that must eventually reach a configuration where no further reactions can occur, regardless of the sequence of reactions executed. This restriction is nontrivial because the techniques of [5, 7] rely on reversible reactions (leading to unbounded executions) catalyzed by species we expect to be depleted once a computational step has terminated. This trick seems to add computational power to our system by undoing certain reactions as long as a specific species is present. Consider the following CRN computing $f(x_{1},x_{2},x_{3})=\min\left(x_{1}-x_{2},x_{3}\right)$. The input values $x_{i}$ are given by the counts of $X_{i}$, and the output by the count of $Z$ molecules in the stable state:
@@
Reactions (1) and (2) compute $x_{1}-x_{2}$, storing the result in the count of $Y$. Next, reaction (3) can be applied exactly $\min(y,x_{3})$ times. But since the order of reactions is a stochastic process, we might consume copies of $Y$ in (3), before all of $x_{2}$ is subtracted from it. Therefore, we add reaction (4), using $X_{2}$ as a catalyst to undo reaction (3) as long as copies of $X_{2}$ are present, indicating that the first step of computation has not terminated. However, this means the above CRN does not have bounded exectutions, since reactions (3) and (4) can be alternated in an infinite execution. A similar technique is used in [5], where semilinear sets are understood as a finite union of linear sets, shown to be computable in parallel by CRNs. A reversible, catalyzed reaction finally converts the output of one of the CRNs to the global output. Among other questions, we explore how the constructions of [5] and [7] can be modified to provide equal computational power while guaranteeing bounded execution.
The paper is organized as follows. Section 3 defines execution boundedness (Definition 3.1). We introduce alternative characterizations of the class for use in later proofs, such as the lack of self-covering execution paths. Section 4 and 5 contain the main positive results of the paper and provide the concrete constructions used to decide semilinear sets and functions using execution bounded CRNs whose initial configurations contain a single leader. Section 6 discusses the limitations of execution bounded CRNs, introducing the concept of a “linear potential function” as a core characterization of these systems. We demonstrate that entirely execution bounded CRNs that are leaderless and non-collapsing (such as all population protocols), can only stably decide trivial semilinear predicates: the eventually constant predicates (Definition 6.11).
@@ 2 Preliminaries
We use established notation from [5, 7] and stable computation definitions from [3] for (discrete) chemical reaction networks.
@@ 2.1 Notation
Let $\mathbb{N}$ denote the nonnegative integers. For any finite set $\Lambda$, we write $\mathbb{N}^{\Lambda}$ to mean the set of functions $f:\Lambda\rightarrow\mathbb{N}$. Equivalently, $\mathbb{N}^{\Lambda}$ can be interpreted as the set of vectors indexed by the elements of $\Lambda$, and so $\mathbf{c}\in\mathbb{N}^{\Lambda}$ specifies nonnegative integer counts for all elements of $\Lambda$. $\mathbf{c}(i)$ denotes the $i$-th coordinate of $\mathbf{c}$, and if $\mathbf{c}$ is indexed by elements of $\Lambda$, then $\mathbf{c}(Y)$ denotes the count of species $Y\in\Lambda$. We sometimes use multiset notation for such vectors, e.g., $\{A,3C\}$ for the vector $(1,0,3)$, assuming there are three species $A,B,C$. If $\Sigma\subseteq\Lambda$, then $\mathbf{i}\upharpoonright\Sigma$ denotes restriction of $\mathbf{i}$ to $\Sigma$.
@@ 2.2 Chemical Reaction Networks
A chemical reaction network (CRN) is a pair $\mathcal{C}=(\Lambda,R)$, where $\Lambda$ is a finite set of chemical species, and $R$ is a finite set of reactions over $\Lambda$, where each reaction is a pair $(\mathbf{r},\mathbf{p})\in\mathbb{N}^{\Lambda}\times\mathbb{N}^{\Lambda}$ indicating the reactants $\mathbf{r}$ and products $\mathbf{p}$. A population protocol [1] is a CRN in which all reactions $(\mathbf{r},\mathbf{p})$ obey $\|\mathbf{r}\|=\|\mathbf{p}\|=2$. (Note that CRNs, including population protocols, do not assume any underlying “communication graph” and model a well-mixed system in which each equal-sized of molecules is as likely to collide and react as any other.) We write reactions such as $A+2B\mathop{\rightarrow}\limits A+3C$ to represent the reaction ($(\{A,2B\},\{A,3C\})$. A configuration $\mathbf{c}\in\mathbb{N}^{\Lambda}$ of a CRN assigns integer counts to every species $S\in\Lambda$. When convenient, we use the notation $\left\{n_{1}S_{1},n_{2}S_{2},\ldots,n_{k}S_{k}\right\}$ to describe a configuration $\mathbf{c}$ with $n_{i}\in\mathbb{N}$ copies of species $S_{i}$, i.e., $\mathbf{c}(S_{i})=n_{i}$, and any species that is not listed is assumed to have a zero count. If some configuration $\mathbf{c}$ is understood from context, for a species $S$, we write #$\#S$ to denote $\mathbf{c}(S).$ A reaction $(\mathbf{r},\mathbf{p})$ is said to be applicable in configuration $\mathbf{c}$ if $\mathbf{r}\leqq\mathbf{c}$. If the reaction $(\mathbf{r},\mathbf{p})$ is applicable, applying it results in configuration $\mathbf{c}^{\prime}=\mathbf{c}-\mathbf{r}+\mathbf{p}$, and we write $\mathbf{c}\rightarrow\mathbf{c}^{\prime}$.
@@
Let $\mathbf{u}\in\mathbb{N}^{R}$. Then the vector $\mathbf{M}\mathbf{u}\in\mathbb{Z}^{\Lambda}$ represents the change in species counts that results from applying reactions by amounts described in $\mathbf{u}$. In the above example, if $\mathbf{u}=(2,1)$, then $\mathbf{M}\mathbf{u}=(-1,0,4)$, meaning that executing the first reaction twice ($\mathbf{u}(1)=2$) and the second reaction once ($\mathbf{u}(2)=1$) causes $S_{1}$ to decrease by 1, $S_{2}$ to stay the same, and $S_{3}$ to increase by 4.
@@ 2.3 Stable computation with CRNs
To capture the result of computations done by a CRN, we generalize the definitions to include information about how to interpret the final configuration after letting the CRN run until the result cannot change anymore (characterized below as stable computation). Computation primarily involves two classes of functions: 1. evaluating predicates $\phi:\mathbb{N}^{k}\to\{0,1\}$ to determine properties of the input, and 2. executing general functions that map an input configuration to an output, denoted as $f:\mathbb{N}^{k}\rightarrow\mathbb{N}$.
The definitions below reference input species $\Sigma\subseteq\Lambda$ and an initial context $\mathbf{s}\in\mathbb{N}^{\Lambda\setminus\Sigma}$. If $\mathbf{s}=\mathbf{0}$ we say that CRN is leaderless. The initial context may be any constant multiset of species, though in practice it tends to be a single “leader” molecule. Furthermore, other initial contexts such as $\{2A,3B\}$ could be produced from a single leader $L$ via a reaction $L\to 2A+3B$, so we may assume without loss of generality that the initial context, if it is nonzero, is simply a single leader. In both cases, we say $\mathbf{i}\in\mathbb{N}^{\Lambda}$ is a valid initial configuration if $\mathbf{i}=\mathbf{s}+\mathbf{x}$, where $\mathbf{x}(S)=0$ for all $S\in\Lambda\setminus\Sigma$; i.e., $\mathbf{i}$ is the initial context plus only input species.
@@
A chemical reaction decider (CRD) is a tuple $\mathcal{D}=(\Lambda,R,\Sigma,\Upsilon_{1},\Upsilon_{0},\mathbf{s})$, where $(\Lambda,R)$ is a CRN, $\Sigma\subseteq\Lambda$ is the set of input species, $\Upsilon_{1}\subseteq\Lambda$ is the set of yes voters, and $\Upsilon_{0}\subseteq\Lambda$ is the set of no voters, such that $\Upsilon_{1}\cap\Upsilon_{0}=\emptyset$, and $\mathbf{s}\in\mathbb{N}^{\Lambda\backslash\Sigma}$ is the initial context. If $\Upsilon_{1}\cup\Upsilon_{0}=\Lambda$, we say the CRD is all-voting. We define a global output partial function $\Phi:\mathbb{N}^{\Lambda}\dashrightarrow\{0,1\}$ as follows. $\Phi(\mathbf{c})$ is undefined if either $\mathbf{c}=\mathbf{0}$, or if there exist $S_{0}\in\Upsilon_{0}$ and $S_{1}\in\Upsilon_{1}$ such that $\mathbf{c}\left(S_{0}\right)>0$ and $\mathbf{c}\left(S_{1}\right)>0$. In other words, we require a unanimous vote as our output. We say $\mathbf{c}$ is stable if, for all $\mathbf{c}^{\prime}$ such that $\mathbf{c}\Rightarrow\mathbf{c}^{\prime}$, $\Phi(\mathbf{c})=\Phi(\mathbf{c}^{\prime}).$ We say a CRD $\mathcal{D}$ stably decides the predicate $\psi:\mathbb{N}^{\Sigma}\rightarrow\{0,1\}$ if, for any valid initial configuration $\mathbf{i}\in\mathbb{N}^{\Lambda}$, letting $\mathbf{i}_{0}=\mathbf{i}\upharpoonright\Sigma$, for all configurations $\mathbf{c}\in\mathbb{N}^{\Lambda},\mathbf{i}\Rightarrow\mathbf{c}$ implies $\mathbf{c}\Rightarrow\mathbf{c}^{\prime}$ such that $\mathbf{c}^{\prime}$ is stable and $\Phi\left(\mathbf{c}^{\prime}\right)=\psi\left(\mathbf{i}_{0}\right)$. We associate to a predicate $\psi$ the set $A=\psi^{-1}(1)$ of inputs on which $\psi$ outputs 1, so we can equivalently say the CRD stably decides the set $A.$
@@
A chemical reaction computer $(CRC)$ is a tuple $\mathcal{C}=(\Lambda,R,\Sigma,Y,\mathbf{s})$, where $(\Lambda,R)$ is a CRN, $\Sigma\subset\Lambda$ is the set of input species, $Y\in\Lambda\backslash\Sigma$ is the output species, and $\mathbf{s}\in\mathbb{N}^{\Lambda\backslash\Sigma}$ is the initial context. A configuration $\mathbf{o}\in\mathbb{N}^{\Lambda}$ is stable if, for every $\mathbf{c}$ such that $\mathbf{o}\Rightarrow\mathbf{c},\mathbf{o}(Y)=\mathbf{c}(Y)$, i.e. the output can never change again. We say that $\mathcal{C}$ stably computes a function $f:\mathbb{N}^{k}\rightarrow\mathbb{N}$ if for any valid initial configuration $\mathbf{i}\in\mathbb{N}^{\Sigma}$ and any $\mathbf{c}\in\mathbb{N}^{\Lambda},\mathbf{i}\Rightarrow\mathbf{c}$ implies $\mathbf{c}\Rightarrow\mathbf{o}$ such that $\mathbf{o}$ is stable and $f(\mathbf{i}\upharpoonright\Sigma)=\mathbf{o}(Y)$.